For the reaction of crystal violet with NaOH(aq), the measured rate of reaction is 1.27 x 10-6 mol L-1 s-1 when the concentration of crystal violet cation is 4.13 x 10-5 mol/L. Estimate how long it will take for the concentration of crystal violet to drop from 4.30 x 10-5 mol/L to 3.96 x 10-5 mol/L.


1
Expert's answer
2021-01-18T01:54:38-0500

Knowing that the reaction is first-ordered:

v=k[A]

k=v/[A]=1.27×10-6/4.13×10-5=0.03(s-1)

ln[A]1=-kt+ln[A]2

t=(ln[A]2-ln[A]1)/k

t=(ln3.96×10-5-ln4.30×10-5)/0.03= (10.14-10.05)/0.03=0.09/0.03=3(s)